# RICE: Breaking Through the Training Bottlenecks of Reinforcement Learning with Explanation

### v1

@@ C.4 Evaluation Results of MuJoCo Games with Sparse Rewards
Results of Sparse HalfCheetah.   First, we compare our refining method with other baseline methods (i.e., PPO fine-tuning, StateMask-R, and JSRL). Figure 9 shows that our refining method is able to help the DRL agent break through the bottleneck with the highest efficiency compared with other baseline refining methods. Additionally, by replacing our explanation method with a random explanation, we observe that the refining performance is getting worse.
Figure 9: Agent Refining Performance in the SparseHalfCheetah Games. For the left figure, we fix the explanation method to our method (mask network) if needed while varying refining methods. For the right figure, we fix the refining method to our method while varying the explanation methods.
@@ D.1 Design Intuitions
  Original agent w/o refine From default initial S {$A_{4}$: 4,914, $A_{9}$: 5} 33.8%
@@ D.2 Discovery of Design Flaws
Second, we find that the intermediate rewards exhibit unusually high sparsity, i.e., many intermediate rewards tend to have a value close to zero, which poses a significant challenge for the PPO algorithm to learn a good policy based on such intermediate rewards. Agents refined with these methods can be easily trapped in local minima. This also explains why AIRS (which uses a reward approximation model) is struggling to provide high-fidelity explanations.
@@
Appendix E Refining an SAC Agent
We claim that our framework is general to refine pre-trained agents that were not trained by PPO algorithms. Here we do experiments on refining an SAC agent in the Hopper game. We obtain a pre-trained SAC agent with $10^{6}$ training steps. However, the pre-trained SAC agent is still far from being optimal. Therefore, it is reasonable for the developer to consider switching to a new DRL algorithm. We utilize the GAIL implementation provided by the imitation library (Gleave et al., 2022) and obtain a new agent that can be trained by PPO algorithm. Figure 3 shows the comparison between using RICE and other baselines to refine the agents. We observe that fine-tuning the DRL agent with SAC algorithm still suffers from the training bottleneck while switching to the PPO algorithm provides an opportunity to break through the bottleneck. Experiments show that our refining scheme is able to help the DRL agent break through the bottleneck and reach a satisfactory performance at the fastest pace. We provide the refining curves when varying hyper-parameters $p$ and $\lambda$ in Figure 13.

### v3

@@ C.4 Evaluation Results of MuJoCo Games with Sparse Rewards
Results of SparseWalker2d.   First, we compare our refining method with other baseline methods (i.e., PPO fine-tuning, StateMask-R, and JSRL) in the SparseWalker2d game. Figure 10 shows that our refining method is able to help the DRL agent break through the bottleneck with the highest efficiency compared with other baseline refining methods. Additionally, by replacing our explanation method with a random explanation, we observe that the refining performance is getting worse.
Figure 10: Agent Refining Performance in the SparseWalker2d Games. For the left figure, we fix the explanation method to our method (mask network) if needed while varying refining methods. For the right figure, we fix the refining method to our method while varying the explanation methods.
@@
C.5 Qualitative Analysis
Figure 14: (a): In the original trajectory, the target agent (the green car) eventually collides with the pink car, which is an undesired outcome. Each time step is marked with a different color: “yellow” indicates the least important step and “red” represents the most important step. (b): We highlight the critical states identified by our explanation method and the corresponding outcome after refining. Using our explanation method, the target agent (the green car) successfully avoids collision.
We do a qualitative analysis of the autonomous driving case to understand how RICE impacts agent behavior and performance. We visualize the agent’s behavior before and after refining the agent. Figure 14(a) shows a trajectory wherein the target agent (depicted by the green car) fails to reach its destination due to a collision with a pink car on the road. Given the undesired outcome, we use our method to identify the critical steps that contribute to the final (undesired) outcome. The important steps are highlighted in red color. Our method identifies the important step as the one when the green car switches across two lanes into the lane of the pink car. The critical state is reasonable because this early step allows the green car to switch lanes to avoid the collision. Based on the provided explanation, we apply our refining method to improve the target agent. The trajectory after refining is shown in Figure 14(b). It shows that after refining, the refined agent (the green car) successfully identifies an alternative path to reach the destination while avoiding collision.
@@ D.1 Design Intuitions
  Original agent w/o refinement From default initial S {$A_{4}$: 4,914, $A_{9}$: 5} 33.8%
@@ D.2 Discovery of Design Flaws
Second, we find that the intermediate rewards exhibit unusually high sparsity, i.e., many intermediate rewards tend to have a value close to zero, which poses a significant challenge for the PPO algorithm to learn a good policy based on such intermediate rewards. Agents refined with these methods can be easily trapped in local minima.
@@
Appendix E Limitation
We use the continuous “Mountain Car” environment (mou, ) as a negative control task to illustrate a scenario where RICE does not work well. In this “extreme” case, Assumption 3.2 is completely broken since the state coverage of the pre-trained agent is limited to a small range around the initial point. In this experiment, we train a target agent using Proximal Policy Optimization (PPO) for 1 million steps. The results show that the policy performance remained poor, with the agent frequently getting trapped at the starting point of the environment. In such cases where the original policy fails to learn an effective strategy, the role of explanations becomes highly limited. Since RICE relies on the identified critical states to enhance the policy, if the policy itself is extremely weak (i.e., not satisfying Assumption 3.2), then the explanations will not be meaningful, which further huts the refinement. In the case of the Mountain Car experiment, RICE essentially reduces to being equivalent to Random Network Distillation (RND) due to the lack of meaningful explanation. We show the result when refining the pre-trained agent using our method and RND in Figure 15.
Figure 15: Refining performance with our method and RND method in MountainCarContinuous-v0 game. The state coverage of the pre-trained policy is limited to a small range around the initial point.